OpenStudy (anonymous):

Solve the problem using the Zero Product Property .. -x(x-2)=0.. Please answer :(

OpenStudy (anonymous):

Everywhere there is an x there is an answer. This is based on the premise that if you multiply two numbers together and the answer is 0 then one of them had to be 0. so you have two numbers here -x and (x - 2) so either -x = 0 or x - 2 = 0 -x = 0 x = 0 x - 2 = 0 x = 2 so x = 0 or x = 2

OpenStudy (anonymous):

Thanks, So much (: